please help. Thank you so much!!(The following information applies to the questions displayed below.) The Gilster Company, a machine tooling firm, has severaC-1. The sales policy at the St. Cloud plant dictates that job bids be calculated by adding 18 percent to total manufacturingc-1. The sales policy at the St. Cloud plant dictates that job bids be calculated by adding 18 percent to total manufacturingd. Compute the under-or overapplied overhead for the St. Cloud plant for the year. (Round your intermediate calculations to 2f. Would your response to part e change if the St. Cloud plant could use the facilities necessary to produce parts for job no
Answer #1

Answer c 1

Bid price = 6.47 + [18%*6.47]  = 7.63 or $8

Overhead rate = (250000+133000+ 288000) / (46000+15000) = 11

Manufacturing costs for Job 110 per unit = 97000 /15000 = 6.47

Answer c 2

Bid price = 4.84 + [18%*4.84]  = 5.71 or $6

Total Manufacturing costs for Job 110 = 19000 + (33000+7600) + (3.73*(120+1200) + 7*120 + 6*1200) = 72563.6

Manufacturing costs for Job 110 per unit = 72563.6 / 15000 = 4.84

Answer c 3

Overhead rate by machine Hours

Answer d

Overhead applied = 3.73*(19500+50000) + 7*19500 + 6*50000

Overhead applied = $ 695735

Actual Overhead = 232000 + 132000+303000

Actual Overhead = 667000

Overapplied Overhead = 695735 - 667000

Overapplied Overhead = $ 28735

answer f is dependent on answer e. No informtaion is provided on part e